lesly327:

An old-fashioned Chinese restaurant offers a family dinner where you get to choose one dish from “column A” (which has 8 dishes), one dish from “column B” (which has 10 dishes) and one dish from “column C” (which has 5 dishes). How many different family dinners can be chosen? There are __________ possible dinners.

AP:

\[A \times B \times C = 8 \times 10 \times 5\]